solve the equation.
3/y=y/5y-12

Found 3 solutions by checkley77, richwmiller, user_dude2008:
Answer by checkley77(12569)

3/y=y/(5y-12) cross multiply.
y^2=3(5y-12)
y^2=15y-36
y^2-15y+36=0
(y-12)(y-3)=0
y-12=0
y=12 ans.
y-3=0
y=3 ans.

Answer by richwmiller(9144)

cross multiply (multiply one side by the denominator of the other side.
to get rid of denominators

Answer by user_dude2008(1861)

3/y=y/(5y-12)

3(5y-12)=y^2

y^2-15y+36=0

(y-12)(y-3)=0

Answers: y=12 or y=3